# SplitOut: Out-of-the-Box Training-Hijacking Detection in Split Learning via Outlier Detection

### v2

@@ 3 SplitOut
The core idea behind SplitOut is for the clients to train the entire network for a short duration using a small fraction (such as $1\%$) of their private data. The gradients collected will then be used as training data for an outlier detection method, which we expect will flag the gradients from an honest server as inliers and those from a training-hijacking server as outliers. The method as we will describe below can be implemented by using an out-of-the-box outlier detection method from libraries such as scikit-learn [39] and requires only the addition of small amounts of bookkeeping code on top an existing SplitNN training setup. By additionally integrating a window-based detection on top of an existing outlier detection algorithm, we further improve our true and false positive rates.
@@ 3.1 Local Outlier Factor (LOF)
An outlier detection (OD) method tries to identify data points that are “different” from the “normal,” e.g. to identify potentially dubious financial transactions. Local outlier factor (LOF) [6, 10, 24] is an unsupervised OD method readily implemented in popular libraries such as scikit-learn [39]. It does not make any distribution assumptions and is density-based. Rather than outputting a binary decision, LOF assigns a local outlier factor score to each point, where points in homogeneous clusters have low scores and those with smaller local densities have high scores. The LOF algorithm works as follows:
@@ 3.2 Collecting Training Data for Outlier Detection
3:  $D$: train set size
4:  $n$: no. of neighbors for LOF
@@
10:  for $(x_{i},y_{i})$leftarrow $train_{\text{SIM}}$ do
@@
23:     if LOF$(train_{\text{LOF}},n)$ classifies majority of $pred_{\text{LOF}}$ as outliers then
24:        Return “attack”.
@@
TABLE I: Attack detection statistics using SplitOut, collected over 100 runs of the first epoch of training with FSHA [36], backdoor [65], and SplitSpy [16], FSHA with multitask learning (FSHA-MT, returning the average of honest and adversarial losses) and an honest server. The true positive rate (TPR) corresponds to the rate of success at detecting feature-space alignment. The false positive rate (FPR) corresponds to the share of honest training runs in which an attack is mistakenly reported. The $t$ field denotes the average point of detection as the share of total batches.
@@
MNIST 1.0 0.011 1.0 0.011 1.0 0.011 1.00 0.012 0
F-MNIST 1.0 0.011 1.0 0.011 1.0 0.011 0.98 0.029 0
CIFAR10 1.0 0.014 1.0 0.013 1.0 0.013 0.98 0.026 0.02
CIFAR100 1.0 0.014 1.0 0.017 1.0 0.013 1.00 0.064 0.02
Figure 3: Results obtained by attackers for the MNIST, F-MNIST, CIFAR10, and CIFAR100 datasets with respect to the detection times (as shown in Table I). The first row displays the original images, and the last row displays the results obtained by a FSHA [36] attacker able to run for an arbitrary duration without being detected.
This seemingly contradicts the goal of collaborative ML, which was to outsource computation. However, as we will demonstrate in more detail, clients using a very small share of their local data for the data collection phase still perform accurate detection. Moreover, better data utilization, rather than purely outsourcing computation, might be a stronger reason for using SplitNN in certain scenarios (e.g. financially/technologically capable institutions such as universities or hospitals working on regulated private data). Since to collect training data for LOF we need only one epoch of training with 1% of the whole data, this is still feasible for many scenarios where SplitNN is employed. Recall that in SplitNN, the clients would train several layers (but not the whole neural network) with 100% of their data for multiple epochs, and increasing the number of client-side layers is recommended to make the private data more difficult to reconstruct from the intermediate outputs [14].
@@ 3.3 Detecting the Attack
For the LOF algorithm, clients need to decide on a hyperparameter $k$, the number of neighbors to consider. We have followed the existing work on automatically choosing the value of $k$ [62], and observed that setting it equal to one less than the number of honest gradients collected (i.e. the highest feasible value) consistently achieves the best performance across all our datasets.
@@
It is possible to employ another outlier detection method in SplitOut besides LOF. We also experimented with other method such as one-class SVM [45] but found results employing LOF to be far superior, and hence present SplitOut and results as using LOF.
@@ 4.1 Setup
For our experiments, we use the ResNet architecture [21], trained using Adam optimizer [27] with learning rate of 0.001, on the MNIST [32], Fashion-MNIST [60], and CIFAR10/100 [30] datasets. With a batch size of 64, one epoch is equal to 938 batches for MNIST and F-MNIST, and 782 batches for CIFAR10/100. We implemented SplitOut in Python (v 3.10) using the PyTorch (v 2.1.0) [37] and scikit-learn (v 1.2) [39] libraries.
We test SplitOut’s performance when trained on with different shares of client training data (1%, 10%, 25%, 50%, 75%, 100%) and different decision-window sizes (1, 10). As stated earlier, we set the number of neighbors to the highest possible value (i.e. one less than the number of training points): 937 for MNIST/F-MNIST, and 781 for CIFAR10/100 when whole training data was used. If the batch size is expressed as $N_{B}$, the selected number of neighbors $k$ can be calculated as $(\lfloor N_{B}\rfloor-1)$.
For each configuration, we measure the true and false positive rates by testing the method on 100 randomly-initialized runs each for the honest training scenario including training LOF from scratch, and each attack. A positive (attack detected) decision in a training-hijacking run is a true positive, and in an honest run is a false positive. To make sure we can detect the attack early enough, we start the algorithm after the required number of gradients for the decision window has been obtained.
TABLE II: SSIM [57] values comparing reconstructed images by the attackers with the original private image averaged over 10 runs. The bottom row displays SSIM results from FSHA’s reconstructed images without detection. A more successful reconstruction process is implied by a higher SSIM value.
MNIST F-MNIST CIFAR CIFAR100
FSHA with SplitOut 0.0004 0.0008 0.0002 0.0002
SplitSpy with SplitOut 0.0002 0.0004 0.0013 0.0001
Without detection 0.8902 0.6640 0.5035 0.2078
@@
4.2 Detecting Training-Hijacking
The effectiveness of SplitOut against all split learning training-hijacking attacks proposed so far (FSHA [36], backdoor attack [65], SplitSpy [16]) is displayed in Table I. We take the window size to be 10 and use as training data for LOF only $1\%$ of the clients’ data, which corresponds to e.g. 9 and 7 training iterations in F-/MNIST and CIFAR10/100 datasets, respectively, with a batch size of 64. More comprehensive results covering a wider range of hyperparameters can be found in Appendix A.
In the more simpler MNIST and F-MNIST datasets, SplitOut achieves a true positive rate (TPR) of 1 and a false positive rate (FPR) of 0, meaning it can detect whether there is an attack or not with perfect accuracy. For the more complex CIFAR datasets, $2\%$ of honest runs are classified as malicious, but that also reduces to zero as more data is allocated for LOF training (see Appendix A), resulting in a higher compute load for the clients. Importantly, LOF’s time to detect attacks does not vary significantly among the different training-hijacking attacks presented so far.
Figure 5: Pair-wise average $L_{2}$ distances between five randomly-chosen sets of gradients (corresponding to the first epoch of training) for honest training, FSHA [36] and the backdoor attack [65], each using the CIFAR10 dataset. In each of the plots, $H$ indicates honest gradients and $M$ malicious. Darker colors imply smaller values.
The structural similarity index measure (SSIM) [57] is a metric used to compare the quality of two different images. In our study, we used SSIM to estimate the image quality of the image reconstructed by the attacker, taking the original input image as a reference. For each dataset, SSIM values of 10 reconstructed images were calculated using the original image as reference, and their averages are shown in Table II. As previously shown in Figure 3 and indicated by the low SSIM values in Table II, SplitOut detects different attacks very early, preventing the malicious server from performing successful image reconstruction.
@@ 4.2 Detecting Training-Hijacking
Figure 6: T-SNE [47] dimension reduction comparing honest and malicious (FSHA & the backdoor attack) gradients obtained from a randomly chosen run of the first epoch of training on CIFAR10.
4.3 Adaptive Attacks
@@
To try to overcome a detection mechanism, a training-hijacking attacker can perform multitask learning and combine the feature-space alignment and honest training losses, e.g. by taking their average, and returning the resulting gradients to the client model. To demonstrate this scenario, we consider a FSHA attacker returning the average of the two losses. Table I displays our results against a multitasking FSHA (denoted FSHA-MT) attacker as well across our four benchmark datasets. We can detect all instances of the attack for on MNIST and CIFAR100 datasets, and 98% of the attack instances on F-MNIST and CIFAR10. Furthermore, Figure 4 displays the reconstruction results the attacker achieves under this setting. Compared to the results in Figure 3, the attack is expectedly less effective, since the client model is now updated to optimize the honest training loss as well as the adversarial discriminator loss.
@@
To show that SplitOut can be used cooperatively with the earlier detection work, we also consider the scenario demonstrated in [16], where the training-hijacking attacker has full knowledge of the detection mechanism in [13]. We use our approach in tandem with the method of [13], and consider only the non-randomized batches as inputs to LOF. We use the name SplitSpy for this approach following [16] with the detection results displayed in Table I, where the LOF algorithm detects all instances of the attack similar to the earlier scenarios. This demonstrates that different detection methods can be used together and that even if the attacker circumvents the detection method of [13], the attack can still be detected using SplitOut.
4.4 Avoiding the Curse of Dimensionality
@@
Our method of choice, the local outlier factor technique, is a distance-based method that makes frequent use of nearest-neighbor computations. Since the data we are working with is very high dimensional (e.g. 1728 dimensions for the gradients from the CIFAR tasks), the curse of dimensionality could have an adverse impact and lead to poor performance. More formally, it has been shown in [4] that for all $L_{p}$ norms with $p\geq 1$, as the dimensionality increases, the proportion $D_{\text{max}}/D_{\text{min}}$ of the distance $D_{\text{max}}$ to the farthest point and $D_{\text{min}}$ to the nearest point vanishes, meaning that the concept of a neighborhood basically collapses. Although $L_{p}$ norms with small $p$ values have been shown to be preferable in higher dimensions [1], they are still subject to the curse of dimensionality [67].
@@
Thus, we now analyze the vast amount of data we collected throughout this work to explain the reasons behind the strong performance in a such high-dimensional setting. Figure 5 displays the pair-wise average $L_{2}$ distances between five randomly-chosen sets of gradients from honest training, as well as FSHA [36] and the backdoor attack [65] all corresponding to the first epoch of training. The results are highly informative about the global and local structure of the problem:
@@
The t-SNE [47] plots (Figure 6) displaying a dimension reduction of honest and malicious gradients over one epoch are consistent with the observations above since they also visualize honest gradients packed tightly, and malicious gradients covering a larger space essentially encompassing the honest gradients. This significant difference between honest and malicious neighborhoods is the main factor why a neighborhood-based method such as LOF succeeds in this task.

### v3

@@ 3 SplitOut
The core idea behind SplitOut is for the clients to train the entire network for a short duration using a small fraction (such as $1\%$) of their private data. The gradients collected will then be used as training data for an outlier detection method, which we expect will flag the gradients from an honest server as inliers and those from a training-hijacking server as outliers. The method as we will describe below can be implemented by using an out-of-the-box outlier detection method from libraries such as scikit-learn [scikit-learn] and requires only the addition of small amounts of bookkeeping code on top an existing SplitNN training setup. By additionally integrating a window-based detection on top of an existing outlier detection algorithm, we further improve our true and false positive rates.
@@ 3.1 Local Outlier Factor (LOF)
An outlier detection (OD) method tries to identify data points that are “different" from the “normal," e.g. to identify potentially dubious financial transactions. Local outlier factor (LOF) [breunig2000lof, chen2010comparison, janssens2009outlier] is an unsupervised OD method readily implemented in popular libraries such as scikit-learn [scikit-learn]. It does not make any distribution assumptions and is density-based. Rather than outputting a binary decision, LOF assigns a local outlier factor score to each point, where points in homogeneous clusters have low scores and those with smaller local densities have high scores. The LOF algorithm works as follows:
@@ 3.2 Collecting Training Data for Outlier Detection
3:  $S$: fct. to get subset (e.g. 1%) of the dataset
4:  $k$: no. of neighbors for LOF
@@
10:  for $(x_{i},y_{i})$leftarrow $S(\text{trainset})$ do
@@
23:     if LOF$(train_{\text{LOF}},k)$ classifies majority of $pred_{\text{LOF}}$ as outliers then
24:        Return “attack".
@@
Figure 2: Results obtained by attackers for the MNIST, F-MNIST, CIFAR10, and CIFAR100 datasets with respect to the detection times (as shown in Table 1). The first row displays the original images, and the last row displays the results obtained by a FSHA [pasquini_unleashing_2021] attacker able to run for an arbitrary duration without being detected.
Table 1: Attack detection statistics using SplitOut, collected over 100 runs of the first epoch of training with FSHA [pasquini_unleashing_2021], the backdoor attack [yuHowBackdoorSplit2023], SplitSpy [fuFocusingPinocchioNose2023], FSHA with multitask learning (FSHA-MT, returning the average of honest and adversarial losses), and an honest server.
@@
MNIST 1 .011 1 .011 1 .011 1.00 .012 0
F-MNIST 1 .011 1 .011 1 .011 0.98 .029 0
CIFAR10 1 .013 1 .013 1 .013 0.98 .026 .02
CIFAR100 1 .013 1 .013 1 .013 1.00 .064 .14
This seemingly contradicts the goal of collaborative ML, which was to outsource computation. However, as we will demonstrate in more detail, clients using a very small share of their local data for the data collection phase still perform accurate detection. Moreover, better data utilization, rather than purely outsourcing computation, might be a stronger reason for using SplitNN in certain scenarios (e.g. financially/technologically capable institutions such as universities or hospitals working on regulated private data, as exemplified in [poirot2019split]). Since to collect training data for LOF we need only one epoch of training with 1% of the whole data, this is still feasible for many scenarios where SplitNN is employed. Recall that in SplitNN, the clients would train several layers (but not the whole neural network) with 100% of their data for multiple epochs, and increasing the number of client-side layers is recommended to make the private data more difficult to reconstruct from the intermediate outputs [erdogan2021unsplit].
@@ 3.3 Detecting the Attack
For the LOF algorithm, clients need to decide on a hyperparameter $k$, the number of neighbors to consider. We have followed the existing work on automatically choosing the value of $k$ [xu2019automatic], and observed that setting it equal to one less than the number of honest gradients collected (i.e. the highest feasible value) consistently achieves the best performance across all our datasets.
@@
It is possible to employ another outlier detection method in SplitOut besides LOF. We also experimented with other method such as one-class SVM [scholkopf1999support] but found results employing LOF to be far superior, and hence present SplitOut and results as using LOF. Experimental results with one-class SVM are shown in the Appendix 0.A.
@@ 4 Evaluation & Discussion
To evaluate our method, we attempt to answer the following reserach questions (RQs):
RQ1
How accurately can SplitOut detect training-hijacking attacks of different kinds? (Section 4.2)
RQ2
How accurately can SplitOut detect training-hijacking attacks performed by a multi-tasking attacker learning both the adversarial and honest objectives? (Section 4.3)
RQ3
How accurately can SplitOut detect training-hijacking attacks performed by the adaptive attacker of [fuFocusingPinocchioNose2023]? (Section 4.3)
RQ4
What can we learn about the nature of training-hijacking attacks from the success of SplitOut? (Section 4.4)
Table 2: Detecting FSHA [pasquini_unleashing_2021] using SplitOut for EMNIST-digits, averaged over 50 runs, with window size 10.
Data % TPR t ($\pm$ SE) FPR .10 1 $4.1\cdot 10^{-5}\pm 0$ 0 .25 1 $4.1\cdot 10^{-5}\pm 0$ 0 1 1 $4.1\cdot 10^{-5}\pm 0$ 0
@@ 4.1 Setup
For our experiments, we use the ResNet architecture [he2015deep], trained using Adam optimizer [kingma_adam_2017] with learning rate of 0.001, on the MNIST [lecun2010mnist], Fashion-MNIST [xiao2017/online], CIFAR10/100 [Krizhevsky09learningmultiple] datasets, and EMNIST [cohen2017emnist] dataset. With a batch size of 64, one epoch is equal to 938 batches for MNIST and F-MNIST, 782 batches for CIFAR10/100, and 3750 batches for EMNIST-digits. We implemented SplitOut in Python (v 3.10) using the PyTorch (v 2.1.0) [pytorch] and scikit-learn (v 1.2) [scikit-learn] libraries.
@@
We test SplitOut’s performance when trained on with different shares of client training data (1%, 10%, 25%, 50%, 75%, 100%) and different decision-window sizes (1, 10). As stated earlier, we set the number of neighbors to the highest possible value (i.e. one less than the number of training points): 937 for MNIST/F-MNIST, 781 for CIFAR10/100, and 3749 for EMNIST-digits, when whole training data was used. If the batch size is expressed as $N_{B}$, the selected number of neighbors $k$ can be calculated as $(\lfloor N_{B}\rfloor-1)$.
Since MNIST and CIFAR datasets are close to each other in terms of total number of batches, the performance of SplitOut for EMNIST-digits dataset containing 3750 batches (with batch size 64) are analyzed. Apart from the selected data rates for other datasets, we evaluated the performance of SplitOut for EMNIST-digits dataset with relatively small data rates like 0.25% (1% of MNIST batches, 9 batches in total), and 0.1% (only 3 batches).
Table 3: FSHA detection statistics using SplitOut where client and server have different model architectures, averaged over 50 runs, with LOF data rate 75% and window size 10.
Dataset TPR t ($\pm$ SE) FPR MNIST 1 .$0107\pm.0000$ 0 F-MNIST 1 .$0107\pm.0001$ 0 CIFAR10 1 .$0148\pm.0003$04 CIFAR100 1 .$0197\pm.0038$06
To test the consistency of the results in case the client cannot know the entire network, we also considered scenarios where the client and the honest server models have different architectures. While the client can deploy a model architecture of 4 ResNet blocks with 1,846,656 trainable parameters, SplitOut’s performance was tested in the case where the actual split learning protocol has a model architecture of 6 ResNet blocks with 3,321,984 trainable parameters.
For each configuration, we measure the true and false positive rates by testing the method on 50 or 100 randomly-initialized runs each for the honest training scenario including training LOF from scratch, and each attack. A positive (attack detected) decision in a training-hijacking run is a true positive, and in an honest run is a false positive. To make sure we can detect the attack early enough, we start the algorithm after the required number of gradients for the decision window has been obtained.
4.2 Detecting Training-Hijacking Attacks (RQ1)
The effectiveness of SplitOut against all split learning training-hijacking attacks proposed so far (FSHA [pasquini_unleashing_2021], backdoor attack [yuHowBackdoorSplit2023], SplitSpy [fuFocusingPinocchioNose2023]) is displayed in Table 1. We take the window size to be 10 and use as training data for LOF only $1\%$ of the clients’ data, which corresponds to e.g. 9 and 7 training iterations in F-/MNIST and CIFAR10/100 datasets, respectively, with a batch size of 64. More comprehensive results covering a wider range of hyperparameters can be found in Appendix 0.B.
@@ 4.2 Detecting Training-Hijacking Attacks (RQ1)
Figure 4: T-SNE [van2008visualizing] dimension reduction comparing honest and malicious (FSHA & the backdoor attack) gradients obtained from a randomly chosen run of the first epoch of training on CIFAR10.
In the more simpler MNIST and F-MNIST datasets, SplitOut achieves a true positive rate (TPR) of 1 and a false positive rate (FPR) of 0, meaning it can detect whether there is an attack or not with perfect accuracy. For the more complex CIFAR datasets, $2\%$ $\sim$ $14\%$ of honest runs are classified as malicious, but that also reduces to zero as more data is allocated for LOF training (see Appendix 0.B), resulting in a higher compute load for the clients. In the dataset with a larger number of batches like EMNIST-digits, SplitOut can detect all attacks even with very small data rates such as 0.1%, 0.25%, and can also differentiate non-attack situations. Importantly, LOF’s time to detect attacks does not vary significantly among the different training-hijacking attacks presented so far.
In a scenario where the client and server have different model architectures and the client cannot know the actual architecture exactly, SplitOut can detect all attacks for all 4 datasets. However, due to reasons such as the client’s model being smaller and the number of parameters being less, LOF should be trained with more data (for the rates we used in our experiments, more than 50%) in order for the false positive rate rate to approach 0 in the CIFAR datasets. When evaluating the success of LOF in this scenario, it should be taken into account that the client has a smaller model in our setup.
Table 4: SSIM [wang-ssim] values comparing reconstructed images by the attackers with the original private image averaged over 10 runs. The bottom row displays SSIM results from FSHA’s reconstructed images without detection. A more successful reconstruction process is implied by a higher SSIM value.
MNIST F-MNIST CIFAR CIFAR100 FSHA w/ SplitOut .$0003\pm.0007$$0006\pm.0010$$0006\pm.0008$$0004\pm.0002$ SplitSpy w/ SplitOut .$0001\pm.0004$$0002\pm.0014$$0005\pm.0011$$\pm.0002$0002 Without detection .$8455\pm.0162$$6770\pm.0199$$5189\pm.0317$$2339\pm.0317$
The structural similarity index measure (SSIM) [wang-ssim] is a metric used to compare the quality of two different images. In our study, we used SSIM to estimate the image quality of the image reconstructed by the attacker, taking the original input image as a reference. For each dataset, SSIM values of 10 reconstructed images were calculated using the original image as reference, and their averages are shown in Table 4. As previously shown in Figure 2 and indicated by the low SSIM values in Table 4, SplitOut detects different attacks very early, preventing the malicious server from performing successful image reconstruction.
4.3 Detecting Adaptive Attacks (RQs 2,3)
@@
To try to overcome a detection mechanism, a training-hijacking attacker can perform multitask learning and combine the feature-space alignment and honest training losses, e.g. by taking their average, and returning the resulting gradients to the client model. To demonstrate this scenario, we consider a FSHA attacker returning the average of the two losses. Table 1 displays our results against a multitasking FSHA (denoted FSHA-MT) attacker as well across our four benchmark datasets. We can detect all instances of the attack on MNIST and CIFAR100 datasets, and 98% of the attack instances on F-MNIST and CIFAR10. Furthermore, Figure 3 displays the reconstruction results the attacker achieves under this setting. Compared to the results in Figure 2, the attack is expectedly less effective, since the client model is now updated to optimize the honest training loss as well as the adversarial discriminator loss.
@@
To show that SplitOut can be used cooperatively with the earlier detection work, we also consider the scenario demonstrated in [fuFocusingPinocchioNose2023], where the training-hijacking attacker has full knowledge of the detection mechanism in [splitguard]. We use our approach in tandem with the method of [splitguard], and consider only the non-randomized batches as inputs to LOF. We use the name SplitSpy for this approach following [fuFocusingPinocchioNose2023] with the detection results displayed in Table 1, where the LOF algorithm detects all instances of the attack similar to the earlier scenarios. This demonstrates that different detection methods can be used together and that even if the attacker circumvents the detection method of [splitguard], the attack can still be detected using SplitOut.
4.4 Avoiding the Curse of Dimensionality (RQ4)
@@
LOF is a distance-based method that makes frequent use of nearest-neighbor computations. Since the data we are working with is very high dimensional, the curse of dimensionality could have an adverse impact and lead to poor performance. More formally, it has been shown in [beyer1999nearest] that for all $L_{p}$ norms with $p\geq 1$, as the dimensionality increases, the proportion $(D_{\text{max}}-D_{\text{min}})/D_{\text{min}}$ with $D_{\text{max}}$ the distance to the farthest point and $D_{\text{min}}$ to the nearest point vanishes, meaning that the concept of a neighborhood basically collapses. Although $L_{p}$ norms with small $p$ values have been shown to be preferable in higher dimensions [aggarwalSurprisingBehaviorDistance2001], they are still subject to the curse of dimensionality [zimekSurveyUnsupervisedOutlier2012].
Figure 5: Pair-wise average $L_{2}$ distances between five randomly-chosen sets of gradients (corresponding to the first epoch) for honest training, FSHA [pasquini_unleashing_2021] and the backdoor attack [yuHowBackdoorSplit2023], each using the CIFAR10 dataset. In each of the plots, $H$ indicates honest gradients and $M$ malicious. Darker colors imply smaller values.
@@
In light of this the counter-intuitive accuracy of LoF requires an explanation. We now analyze the vast amount of data we collected throughout this work to explain the reasons behind the strong performance in a such high-dimensional setting. Figure 5 displays the pair-wise average $L_{2}$ distances between five randomly-chosen sets of gradients from honest training, as well as FSHA [pasquini_unleashing_2021] and the backdoor attack [yuHowBackdoorSplit2023] all corresponding to the first epoch of training. The results are highly informative about the global and local structure of the problem:
@@
The t-SNE [van2008visualizing] plots (Figure 4) displaying a dimension reduction of honest and malicious gradients over one epoch are consistent with the observations above since they also visualize honest gradients packed tightly, and malicious gradients covering a larger space essentially encompassing the honest gradients. This significant difference between honest and malicious neighborhoods is the main factor why a neighborhood-based method such as LOF succeeds in this task.